# Point-DeepONet: A Deep Operator Network Integrating PointNet for Nonlinear Analysis of Non-Parametric 3D Geometries and Load Conditions

### v1

@@ 3.4 Effect of Input Configurations on Model Performance
Table 5: Performance comparison of PointNet, DeepONet, and Point-DeepONet models across various input configurations for displacement and von Mises stress metrics
@@
Model Input Parameters Component $u_{x}$ (mm) $u_{y}$ (mm) $u_{z}$ (mm) von Mises stress (MPa) MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ PointNet $X$, $f$, $\mathbf{d}$ Vertical 0.008 0.018 0.950 0.003 0.006 0.920 0.014 0.034 0.951 12.111 18.671 0.860 Horizontal 0.006 0.011 0.982 0.002 0.003 0.876 0.007 0.012 0.982 9.042 14.183 0.904 Diagonal 0.004 0.005 0.884 0.002 0.003 0.960 0.006 0.009 0.968 8.224 12.406 0.844 $X$, $m$, $f$, $\mathbf{d}$ Vertical 0.008 0.018 0.953 0.003 0.006 0.927 0.013 0.034 0.952 11.666 18.291 0.866 Horizontal 0.006 0.010 0.983 0.002 0.003 0.878 0.008 0.012 0.980 9.073 14.208 0.904 Diagonal 0.004 0.005 0.884 0.002 0.003 0.963 0.006 0.008 0.975 7.639 11.786 0.859 DeepONet Branch: $f$, $\mathbf{d}$ Trunk: $X$, $SDF$ Vertical 0.026 0.051 0.625 0.007 0.012 0.709 0.048 0.090 0.653 21.893 31.979 0.590 Horizontal 0.024 0.044 0.692 0.004 0.006 0.503 0.026 0.047 0.700 20.187 29.984 0.570 Diagonal 0.006 0.010 0.578 0.005 0.007 0.752 0.016 0.024 0.773 12.607 18.743 0.644 Branch: $m$, $f$, $\mathbf{d}$ Trunk: $X$ Vertical 0.019 0.037 0.795 0.006 0.010 0.775 0.032 0.061 0.841 19.081 28.328 0.678 Horizontal 0.017 0.033 0.835 0.003 0.005 0.561 0.019 0.035 0.835 17.867 26.777 0.657 Diagonal 0.006 0.010 0.620 0.004 0.007 0.787 0.013 0.020 0.854 11.930 17.973 0.672 Branch: $m$, $f$, $\mathbf{d}$ Trunk: $X$, $SDF$ Vertical 0.026 0.050 0.626 0.007 0.012 0.709 0.048 0.090 0.654 21.842 31.901 0.592 Horizontal 0.024 0.044 0.694 0.004 0.006 0.503 0.026 0.047 0.699 20.091 29.935 0.572 Diagonal 0.006 0.010 0.578 0.005 0.007 0.753 0.016 0.024 0.774 12.555 18.710 0.645 Point-DeepONet (Our proposed) Branch: $f$, $\mathbf{d}$ Trunk: $X$, $SDF$ Vertical 0.007 0.017 0.959 0.003 0.006 0.919 0.032 0.058 0.958 10.538 17.043 0.883 Horizontal 0.005 0.010 0.983 0.002 0.003 0.876 0.006 0.011 0.983 8.078 12.976 0.920 Diagonal 0.003 0.005 0.905 0.002 0.003 0.950 0.005 0.009 0.972 7.180 11.468 0.867 Branch: $m$, $f$, $\mathbf{d}$ Trunk: $X$ Vertical 0.007 0.017 0.958 0.003 0.006 0.920 0.012 0.033 0.982 8.967 14.347 0.902 Horizontal 0.002 0.003 0.855 0.006 0.011 0.982 0.008 0.012 0.980 9.073 14.208 0.904 Diagonal 0.003 0.005 0.911 0.002 0.003 0.953 0.005 0.008 0.978 7.661 12.028 0.853 Branch: $m$, $f$, $\mathbf{d}$ Trunk: $X$, $SDF$ Vertical 0.007 0.016 0.964 0.003 0.006 0.923 0.012 0.030 0.961 10.541 17.148 0.882 Horizontal 0.005 0.010 0.985 0.002 0.003 0.875 0.006 0.011 0.985 7.935 12.684 0.923 Diagonal 0.003 0.005 0.903 0.002 0.003 0.953 0.005 0.008 0.976 7.090 11.208 0.873 1. *Input Parameters: $X$: coordinates $(x,y,z)$; $SDF$: signed distance function; $m$: mass; $f$: force magnitude; $\mathbf{d}$: direction vector $(d_{x},d_{y},d_{z})$.
For the PointNet model, two input configurations were considered. The first configuration included the spatial coordinates $X$, force magnitude $f$, and direction vector $\mathbf{d}$. The second configuration added mass $m$ to these inputs, resulting in the inclusion of $X$, $m$, $f$, and $\mathbf{d}$. The results indicate that incorporating mass in the input slightly improves performance in certain cases. For instance, when comparing the input configuration with mass to the one without, the $R^{2}$ score for $u_{y}$ under the vertical load case increased from 0.920 to 0.927, and the MAE for von Mises stress under the diagonal load case decreased from 8.224 MPa to 7.639 MPa. This suggests that including mass helps the model better capture the physical behavior of the structure.
The DeepONet model was tested with three input configurations. In the first configuration, the branch inputs were force magnitude $f$ and direction vector $\mathbf{d}$, and the trunk inputs were spatial coordinates $X$ and the signed distance function $SDF$. In the second configuration, the branch inputs included mass $m$, force magnitude $f$, and direction vector $\mathbf{d}$, while the trunk inputs were only the spatial coordinates $X$. The third configuration combined all these inputs, with the branch receiving $m$, $f$, and $\mathbf{d}$, and the trunk receiving $X$ and $SDF$. Overall, the DeepONet model exhibited lower predictive performance compared to PointNet and Point-DeepONet across all configurations, with $R^{2}$ scores generally below 0.8. Notably, including the signed distance function in the trunk did not lead to significant improvements in performance. For example, under the vertical load case with $m$, $f$, and $\mathbf{d}$ in the branch and $X$ and $SDF$ in the trunk, the $R^{2}$ score for $u_{x}$ displacement was only 0.626.
For the Point-DeepONet model, we explored three input configurations similar to those used for DeepONet. The first configuration had branch inputs of $f$ and $\mathbf{d}$, and trunk inputs of $X$ and $SDF$. The second configuration included mass $m$ in the branch inputs, resulting in branch inputs of $m$, $f$, and $\mathbf{d}$, with trunk inputs as $X$. The third configuration combined all inputs, with the branch receiving $m$, $f$, and $\mathbf{d}$, and the trunk receiving $X$ and $SDF$. The results demonstrate that Point-DeepONet consistently outperformed both PointNet and DeepONet across all input configurations and loading conditions. Notably, including the signed distance function in the trunk network led to the highest $R^{2}$ scores in several cases. For example, under the vertical load case with $m$, $f$, and $\mathbf{d}$ in the branch and $X$ and $SDF$ in the trunk, Point-DeepONet achieved an $R^{2}$ score of 0.964 for $u_{x}$ displacement, surpassing the corresponding scores of PointNet and DeepONet.
Comparing the input configurations for Point-DeepONet, the use of mass and the signed distance function in combination with other inputs generally resulted in better performance. The model achieved the highest $R^{2}$ scores and lowest errors when both $m$, $f$, and $\mathbf{d}$ were used in the branch and $X$ and $SDF$ in the trunk. This suggests that incorporating additional physical parameters such as mass and geometric information like the SDF enhances the model’s ability to capture the complex elastoplastic behavior of structures.
In terms of predicting von Mises stress, Point-DeepONet with $m$, $f$, and $\mathbf{d}$ in the branch and $X$ and $SDF$ in the trunk achieved the highest $R^{2}$ scores across all load cases, indicating superior performance in capturing stress distributions. For instance, under the horizontal load case, the $R^{2}$ score for von Mises stress was 0.923, higher than the scores achieved by PointNet and DeepONet. This demonstrates the model’s enhanced capability in predicting stress fields when enriched input features are provided.
Overall, the analysis indicates that the choice of input parameters significantly affects model performance. Point-DeepONet shows robustness and superior predictive capabilities across various input configurations, particularly when incorporating mass and the signed distance function. These findings highlight the importance of selecting appropriate input features that capture both the physical properties and geometric characteristics of structures to enhance model accuracy.
@@ 4 Conclusion and Future Work
This study introduced Point-DeepONet, a novel neural network architecture that integrates PointNet into the DeepONet framework to overcome key challenges in nonlinear elastoplastic analyses of non-parametric, three-dimensional geometries. Our results highlight how the method advances beyond existing surrogate modeling approaches by enabling accurate, efficient, and scalable predictions directly on unstructured finite element meshes. More specifically, our findings support the following conclusions, providing quantitative evidence for each of the key advantages outlined in the Introduction:
@@
    Flexible handling of non-parametric 3D geometries: By leveraging PointNet’s capability to process raw point cloud data, Point-DeepONet effectively modelled a broad range of complex geometries without any parameterization or regular grid mapping. With a resampling size of $5{,}000$ points (only 2.5% of the average $\sim$200,000 nodes in the original finite element meshes), our model achieved an average $R^{2}$ of $0.934$, thereby demonstrating the ability to preserve geometric fidelity and local feature details even at significantly reduced input sizes.
@@
    Versatile adaptation to varying load conditions: Unlike Geom-DeepONet [29], which focuses primarily on load magnitude variations, Point-DeepONet seamlessly accommodates changes in both load magnitude and direction. Under horizontal loading, for instance, the model attained $R^{2}$ values as high as $0.985$ for displacement components and $0.923$ for von Mises stress, outperforming both PointNet and DeepONet configurations. These results confirm that our model can reliably adapt to evolving and non-uniform loading scenarios.
@@
    High-fidelity field predictions at the source mesh: Point-DeepONet provides direct predictions of displacement ($u_{x}$, $u_{y}$, $u_{z}$) and von Mises stress on the original finite element mesh. This eliminates the need for interpolation or additional post-processing. For diagonal load cases, the model delivered $R^{2}$ scores around $0.95$ for displacement fields and $0.87$ for stress predictions, ensuring robust high-resolution field outputs that closely match the finite element analysis (FEA) results.
@@
    Reduced computational overhead: Comparisons with conventional nonlinear FEA highlight significant computational savings. While the average FEA runtime exceeded 1,000 seconds per case, Point-DeepONet required only a few seconds for full-mesh inference, offering a more than $4\times 10^{2}$-fold speedup. This remarkable gain in efficiency enables the rapid exploration of diverse design alternatives and load scenarios.
@@
    Accelerated decision-making processes: Owing to its rapid inference and high predictive accuracy, Point-DeepONet facilitates quick evaluations of numerous design iterations. This advantage was evident when scaling up the dataset size: for instance, with 3,000 training samples, our model sustained an average $R^{2}$ of $0.934$ and maintained training times under an hour, significantly expediting the decision-making process in engineering applications.
@@
    Robust scalability with increasing data and complexity: As the number of training samples and the fidelity of geometric detail grew, Point-DeepONet not only improved in accuracy but also maintained manageable training times. Beyond a dataset size of about 261 samples, it consistently surpassed PointNet in both prediction quality and computational efficiency. This scalability ensures the method remains effective as problem sizes, design spaces, and data availability continue to expand.
Building on these accomplishments, future research can extend Point-DeepONet to encompass a wider spectrum of material models, including anisotropic and composite materials, as well as more complex load histories (e.g., dynamic and time-dependent conditions). Integrating physics-informed neural networks (PINNs) could further enhance model robustness by embedding physical constraints directly into the loss function, potentially reducing data requirements and increasing accuracy in sparse data regimes. Additionally, exploring transfer learning techniques may enable the model to adapt to new geometries and load conditions with minimal retraining, further improving computational efficiency.
@@ Acknowledgments
This work was supported by the Ministry of Science and ICT of Korea grant (No.2022-0-00969, No.2022-0-00986, and GTL24031-300) and the Ministry of Trade, Industry & Energy (RS-2024-00410810).

### v2

@@ 3.4 Ablation Study and Architectural Validation
Table 9: Performance comparison of PointNet, DeepONet, and Point-DeepONet models across various input configurations for displacement and von Mises stress metrics.
@@
Model Input Parameters Component $u_{x}$ ($m$) $u_{y}$ ($m$) $u_{z}$ (mm) von Mises stress (MPa) MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ MAE RMSE $R^{2}$ PointNet $X$, $f$, $\mathbf{d}$ Vertical 0.008 0.018 0.950 0.003 0.006 0.920 0.014 0.034 0.951 12.111 18.671 0.860 Horizontal 0.006 0.011 0.982 0.002 0.003 0.876 0.007 0.012 0.982 9.042 14.183 0.904 Diagonal 0.004 0.005 0.884 0.002 0.003 0.960 0.006 0.009 0.968 8.224 12.406 0.844 $X$, $m$, $f$, $\mathbf{d}$ Vertical 0.008 0.018 0.953 0.003 0.006 0.927 0.013 0.034 0.952 11.666 18.291 0.866 Horizontal 0.006 0.010 0.983 0.002 0.003 0.878 0.008 0.012 0.980 9.073 14.208 0.904 Diagonal 0.004 0.005 0.884 0.002 0.003 0.963 0.006 0.008 0.975 7.639 11.786 0.859 DeepONet Branch: $f$, $\mathbf{d}$ Trunk: $X$, SDFSDF Vertical 0.026 0.051 0.625 0.007 0.012 0.709 0.048 0.090 0.653 21.893 31.979 0.590 Horizontal 0.024 0.044 0.692 0.004 0.006 0.503 0.026 0.047 0.700 20.187 29.984 0.570 Diagonal 0.006 0.010 0.578 0.005 0.007 0.752 0.016 0.024 0.773 12.607 18.743 0.644 Branch: $m$, $f$, $\mathbf{d}$ Trunk: $X$ Vertical 0.019 0.037 0.795 0.006 0.010 0.775 0.032 0.061 0.841 19.081 28.328 0.678 Horizontal 0.017 0.033 0.835 0.003 0.005 0.561 0.019 0.035 0.835 17.867 26.777 0.657 Diagonal 0.006 0.010 0.620 0.004 0.007 0.787 0.013 0.020 0.854 11.930 17.973 0.672 Branch: $m$, $f$, $\mathbf{d}$ Trunk: $X$, SDFSDF Vertical 0.026 0.050 0.626 0.007 0.012 0.709 0.048 0.090 0.654 21.842 31.901 0.592 Horizontal 0.024 0.044 0.694 0.004 0.006 0.503 0.026 0.047 0.699 20.091 29.935 0.572 Diagonal 0.006 0.010 0.578 0.005 0.007 0.753 0.016 0.024 0.774 12.555 18.710 0.645 Point-DeepONet (Our proposed) Branch: $f$, $\mathbf{d}$ Trunk: $X$, SDFSDF Vertical 0.007 0.017 0.959 0.003 0.006 0.919 0.032 0.058 0.958 10.538 17.043 0.883 Horizontal 0.005 0.010 0.983 0.002 0.003 0.876 0.006 0.011 0.983 8.078 12.976 0.920 Diagonal 0.003 0.005 0.905 0.002 0.003 0.950 0.005 0.009 0.972 7.180 11.468 0.867 Branch: $m$, $f$, $\mathbf{d}$ Trunk: $X$ Vertical 0.007 0.017 0.958 0.003 0.006 0.920 0.012 0.033 0.982 8.967 14.347 0.902 Horizontal 0.002 0.003 0.855 0.006 0.011 0.982 0.008 0.012 0.980 9.073 14.208 0.904 Diagonal 0.003 0.005 0.911 0.002 0.003 0.953 0.005 0.008 0.977 7.661 12.028 0.853 Branch: $m$, $f$, $\mathbf{d}$ Trunk: $X$, SDFSDF Vertical 0.007 0.016 0.964 0.003 0.006 0.923 0.012 0.030 0.961 10.541 17.148 0.882 Horizontal 0.005 0.010 0.985 0.002 0.003 0.875 0.006 0.011 0.985 7.935 12.684 0.923 Diagonal 0.003 0.005 0.903 0.002 0.003 0.953 0.005 0.008 0.978 7.090 11.208 0.873 1. *Input Parameters: $X$: coordinates $(x,y,z)$; SDFSDF: signed distance function; $m$: mass; $f$: force magnitude; $\mathbf{d}$: direction vector $(d_{x},d_{y},d_{z})$.
3.5 Validation of Generalization to Unseen Load Conditions
To rigorously evaluate the model’s ability to generalize across both load directions and magnitudes, we conducted a targeted validation study. A new test set was generated using 20 distinct geometries, a subset of which are shown with IDs in Figure 4. For these geometries, we created 50 test cases where load directions were sampled uniformly at random from 0 to 90 degrees, and magnitudes were sampled from a range of 35-45 kN, ensuring the conditions were unseen during training. The Point-DeepONet model, trained on the original large-scale dataset, was then evaluated on these 50 new cases without any fine-tuning. The overall performance, summarized in Table 10, shows only a minor degradation compared to the original validation set, confirming robust generalization.
To further dissect the model’s performance and identify its limitations, we analyzed the error trends from these 50 random test cases, as shown in Figure 21. The scatter plot of error versus load angle (Figure 21a) reveals no discernible correlation, suggesting that the model’s accuracy is largely independent of the specific load direction. This provides strong evidence that the model has successfully generalized across the entire directional space. Similarly, the relationship between error and load magnitude (Figure 21b) appears random, indicating uniform performance across the tested magnitude range.
In contrast, the box plot of errors per geometry (Figure 21c) highlights that certain geometries (e.g., ID 7, 18) consistently yield higher prediction errors regardless of the load condition. A qualitative review of these ”difficult” geometries, by referencing them back to Figure 4, reveals they possess challenging features. For example, Geometry 7 has a deep, sharp U-shaped groove known to cause stress concentration, while Geometry 18 features complex intersections of thin reinforcing members. This finding points to a key limitation of our current approach: while robust to load variations, its accuracy can be sensitive to specific geometric complexities where high-frequency stress concentrations are likely to occur. This analysis provides valuable insight for future model improvements.
Table 10: Performance on the validation set (Fixed Directions) vs. the new test set (Random Directions).
Metric Validation Set (Fixed Dir.) Random-Direction Test Set
Avg. $R^{2}$ (Displacement) 0.970 0.938
Avg. $R^{2}$ (von Mises) 0.898 0.847
Figure 21: Detailed analysis of error trends on the random load test set. (a) MAE vs. load angle shows no clear correlation, indicating good directional generalization. (b) MAE vs. load magnitude is also randomly distributed. (c) Box plot of MAE per geometry reveals that specific geometries (e.g., ID 7, 18) consistently produce higher errors, highlighting a limitation related to geometric complexity.
@@ 4 Conclusion and Future Work
This study introduced Point-DeepONet, a novel neural network architecture that integrates PointNet into the DeepONet framework to overcome key challenges in nonlinear elastoplastic analyses of non-parametric, three-dimensional geometries. Our results highlight how the method advances beyond existing surrogate modeling approaches by enabling accurate, efficient, and scalable predictions directly on unstructured finite element meshes. More specifically, our findings support the following conclusions:
@@
    Flexible handling of non-parametric 3D geometries: By leveraging PointNet’s capability to process raw point cloud data, Point-DeepONet effectively modelled a broad range of complex geometries without any parameterization or regular grid mapping. With a resampling size of 5,0005{,}000 points (only 2.5% of the average mesh), our model achieved an average $R^{2}$ of $0.934$, thereby demonstrating the ability to preserve geometric fidelity even at significantly reduced input sizes.
@@
    Versatile adaptation to varying load conditions: Unlike methods limited to load magnitude variations, Point-DeepONet seamlessly accommodates changes in both load magnitude and direction. Our extensive validation, including tests on unseen interpolated and completely random load directions, confirmed the model’s strong generalization capability, maintaining high accuracy without retraining.
@@
    High-fidelity field predictions at the source mesh: Point-DeepONet provides direct predictions of displacement and von Mises stress on the original finite element mesh, eliminating the need for interpolation. The model delivered robust high-resolution field outputs that closely match the FEA results across a wide range of conditions.
@@
    Reduced computational overhead: Comparisons with conventional nonlinear FEA highlight significant computational savings. While the average FEA runtime exceeded 19 minutes per case, Point-DeepONet required only a few seconds for full-mesh inference, offering a more than $4\times 10^{2}$-fold speedup. Our quantitative analysis also confirmed that the one-time preprocessing cost for SDF generation is negligible, preserving the model’s overall efficiency for iterative applications.
@@
    Accelerated decision-making processes: Owing to its rapid inference and high predictive accuracy, Point-DeepONet facilitates quick evaluations of numerous design iterations. This advantage was evident when scaling up the dataset size, significantly expediting the decision-making process in engineering applications.
@@
    Robust scalability with increasing data and complexity: The model’s performance was validated on a comprehensive dataset of 3,000 simulations. The results from this dataset, combined with ablation studies and targeted generalization tests, confirm that the method is robust and that its key architectural components are essential for achieving high accuracy.
Building on these accomplishments, future research can extend Point-DeepONet to encompass a wider spectrum of material models, including anisotropic and composite materials. While the current study focuses on static loads, the architectural design naturally extends to time-dependent conditions. The branch network can incorporate temporal information through recurrent layers (LSTM/GRU), sequential operator frameworks such as s-DeepONet [11], or attention mechanisms, enabling transient response predictions under seismic, cyclic, or impact loading. Integrating physics-informed neural networks (PINNs) could further enhance model robustness by embedding physical constraints directly into the loss function, potentially reducing data requirements and increasing accuracy in sparse data regimes. Furthermore, our error analysis indicates that model accuracy is sensitive to certain geometric complexities. Future work could address this limitation by incorporating adaptive point sampling strategies that focus on regions with high geometric gradients or by exploring more advanced geometric feature extractors capable of capturing fine-grained local details more effectively. Additionally, exploring transfer learning techniques may enable the model to adapt to new geometries and load conditions with minimal retraining, further improving computational efficiency.
@@ Acknowledgments
This work was supported by the Ministry of Science and ICT of Koreagrant (GTL24031-000, N10250154) and the Ministry of Trade, Industry& Energy (RS-2024-00410810, RS-2025-02317327)